OpenStudy (aaronandyson):

how much heat energy is released when 5g of water at 20 degrees changes into ice at 0 degrees C? take specific heat capacity of water = 4.2j g k and specific latent of fusion of ice as 336

OpenStudy (aaronandyson):

@Michele_Laino

OpenStudy (michele_laino):

we have to apply this formula: \[Q = Lm + cm\Delta \theta \] since, we have subtract heat in order to go from water at 20 degrees to water at 0 degrees, plus we have to subtract heat to go from water at 0 degrees to ice at zero degrees

OpenStudy (aaronandyson):

What do you mean..??

OpenStudy (michele_laino):

after a substitution, we get: \[\Large Q = Lm + cm\Delta \theta = \left( {336 \cdot 5} \right) + \left( {4.2 \cdot 5 \cdot 20} \right) = ...Joules\]

OpenStudy (aaronandyson):

2100 J?

OpenStudy (michele_laino):

here the physical system, namely the sample of water releases energy to the surrounding environment. Such energy (heat) is composed by 2 parts or 2 amounts: 1) the energy released during the temperature decresing, from 20 to 0 degrees 2) the energy released by the so formed sample of water at 0 degrees to form a corresponding sample of ice at zero degrees

OpenStudy (michele_laino):

I got this: the sample of water releases \(2130\) Joules to the surrounding environment